OpenStudy (astrophysics):

Prove each identity, assuming that S and E satisfy the conditions of Divergence Theorem and the scalar functions and components of the vector fields have continuous second order partial derivatives. (Note: I haven't gone through all of them yet, but I have found the solutions to the proofs as I'm trying to figure out some myself still, but if you would like to take a attempt first go ahead, and I will post the solution afterwards).

OpenStudy (bohotness):

Curl (aka rotor) of any 3D vector field is divergenceless: div (curl F(r)) = by components: div [∂Fz/∂y - ∂Fy/∂z, ∂Fx/∂z - ∂Fz/∂x, ∂Fy/∂x - ∂Fx/∂y] = ∂/∂x(∂Fz/∂y - ∂Fy/∂z) + ∂/∂y(∂Fx/∂z - ∂Fz/∂x) + ∂/∂z(∂Fy/∂x - ∂Fx/∂y) = ∂²Fz/∂x∂y - ∂²Fy/∂x∂z + ∂²Fx/∂y∂z - ∂²Fz/∂y∂x + ∂²Fy/∂z∂x - ∂²Fx/∂z∂y = = zero According to Gauss theorem surface flux double integral of curl F dS = is equal to volume integral of divergence = triple integral div curl F dV = = triple integral of ZERO dV = = zero Source(s): http://en.wikipedia.org/wiki/Curl http://en.wikipedia.org/wiki/Divergence_...

OpenStudy (astrophysics):

\[1. \int\limits \int\limits \textbf{a} \cdotp \textbf{n} dS = 0\] where a is a constant vector. 2. \[V(E) = \frac{ 1 }{ 3 } \int\limits \int\limits_S \textbf{F} \cdotp d\textbf{S}\] where F(x,y,z) = xi++yi+zk (i,j,k are unit vectors) \[3. \int\limits \int\limits _S curl \textbf{F} \cdotp d\textbf{S} = 0 \] \[4. \int\limits \int\limits_S D_n f dS = \int\limits \int\limits \int\limits _e \nabla ^2 f dV\] \[5. \int\limits \int\limits_S (f \nabla g) \cdotp \textbf{n} dS = \int\limits \int\limits \int\limits_E (f \nabla ^2 g+ \nabla f \cdotp \nabla g) dV\] \[6. \int\limits \int\limits_S (f \nabla g- g \nabla f) \cdot \textbf{n} dS= \int\limits \int\limits \int\limits ( f \nabla ^2 g - g \nabla ^2 f) dV\] This should be fun, one of them should look familiar to some of you as I've asked it before. Also note solving one may help you with another!

OpenStudy (astrophysics):

@rational @dan815 @kainui @michele_laino @zarkon tag anyone else you think would be interested as well :)

ganeshie8 (ganeshie8):

*

OpenStudy (dan815):